OpenStudy (anonymous):

The sum of a number its square is 30. Find the number.

OpenStudy (lalaly):

let the number be x\[x+x^2=30\]\[x^2+x-30=0\]\[(x+6)(x-5)=0\]\[x=-6 ,5\]

OpenStudy (campbell_st):

so x + x^2 = 30 then the have the quadratic x^2 +x - 30 = 0 (x - 5)(x+6)=0 then x = 5 or -6

OpenStudy (lalaly):

-6 and 5 both work

OpenStudy (campbell_st):

why... (-6)^2 = 36 36 + (-6) = 30.... condition met

OpenStudy (lalaly):

exactly^^